Z varies jointly as x^2 and y, and z = 20 when x = 2 and y = 3. Find z if x = 4 and y = 7/10
Answer by Cromlix(4381)

Hi there,
z varies jointly as x^2 and y
z = kx^2y
k is the constant of proportionality.
To find k use values, z = 20, x = 2 and y = 3
z = kx^2y
20 = k(2)^2 * 3 (* = times)
20 = k12
k = 20/12
k = 1.67 (2 decimal places)
Formula:
z = 1.67 * x^2 * y
Use formula:
z = 1.67x^2y
z = 1.67 * (4)^2 * 7/10
z = 18.70 (2 decimal places)
Hope this helps :-)